# Learning When to Plan: Efficiently Allocating Test-Time Compute for LLM Agents

### v1

@@ B.1 Crafter
Crafter Hafner [2022] is an open-world survival game inspired by Minecraft. The environment is designed to benchmark reinforcement learning agents on tasks requiring generalization and long-term reasoning without extensive prior game knowledge. Crafter presents a procedurally generated 2D world where the agent must gather resources, craft tools, and defend against creatures to survive and unlock achievements. The environment provides observations as a combination of top-down pixel views of the agent’s local surroundings and non-visual data, including the agent’s health, inventory, and currently held item. The agent can perform 17 discrete actions, and its progress is measured by unlocking 22 predefined achievements, each yielding a sparse reward signal.
Our agent interacts with the Crafter environment through the BALROG Paglieri et al. [2025a] framework, which acts as a standardized wrapper. BALROG facilitates the communication between our agent, conceptualized as an underlying Large Language Model (LLM) combined with a specific prompting strategy, and the Crafter game environment. At each timestep, Crafter, via BALROG, relays the current observation to our agent. The agent, which internally maintains a history of past observations and actions, incorporates this new observation into its context. A dedicated prompt builder component within the agent updates this interaction history and formats it into a chat template.
This prompt is then passed to the LLM, which processes the contextual information and generates the subsequent action as a natural language string. BALROG then translates this string into a game-compatible command for Crafter. Further details regarding the specifics of the system prompt, the wrapper’s mechanics, and the handling of action validation can be found in BALROG Paglieri et al. [2025a].
@@ C.1 Supervised Fine-Tuning
For the Supervised Fine-Tuning (SFT) phase, we generated a dataset of 1,024 trajectories using Llama 3.3 70B instruct. To ensure data diversity, we employed 16 distinct planning prompts (detailed in Appendix A), which were sampled uniformly. The planning frequency for trajectories in the dataset was also sampled uniformly from the range [2, 12]. The Llama 3.1 8B instruct model served as the baseline for all SFT experiments, with no Low-Rank Adaptation (LoRA) adapters applied during this stage [Hu et al., 2022]. We used AdamW [Loshchilov and Hutter, 2019] as the optimizer throughout.
We considered two objective functions: direct action prediction and full world modeling. Our findings, presented in Figures 7 and 8, indicate that omitting world modeling yields superior results in terms of both performance and reduced catastrophic forgetting. While we initially tested world modeling as a potential regularization technique to mitigate overfitting, it ultimately proved to be a distractor for the agents. To additionally model the environment dynamics, the weights of the model had to be changed more significantly, which is reflected in the higher KL divergence, and consequently led to increased forgetting and reduced model steerability. It is worth noting that for the agent to plan dynamically during SFT, it must also model the plans themselves, exposing it to a greater number of tokens during this stage. Training was performed using the DeepSpeed ZeRO Stage 3 optimizer [Aminabadi et al., 2022] to effectively manage memory and scale operations across multiple GPUs.
@@ C.2 RLFT
During the Reinforcement Learning Fine-Tuning (RLFT) phase, we applied RSLoRA adapters [Kalajdzievski, 2023] separately to actor and critic models. Data collection involved using vLLM [Kwon et al., 2023] to host model weights and BALROG [Paglieri et al., 2025a] integrated with Ray [Moritz et al., 2018] to process outputs and handle agent-environment interactions efficiently. A custom Proximal Policy Optimization (PPO) trainer [Schulman et al., 2017] was implemented, with model weights broadcasted back to vLLM after each training epoch. The actor model was based on the 8B parameter Llama 3.1 architecture, while the critic model utilized the 1B parameter Llama 3.3 architecture [Grattafiori et al., 2024]. Similar to the SFT phase, the DeepSpeed ZeRO Stage 3 optimizer was employed for memory-efficient and scalable training.
Experiments were conducted using a total of 8 GPUs, allocating 6 GPUs for training and 2 for data collection. Most experiments were executed on a node with 8xH100 GPUs and typically completed within 24–48 hours. Initially, we experimented with reward penalties targeting invalid actions, excessively long responses, and overly frequent planning. However, such explicit reward shaping often resulted in agents refraining from planning entirely. Recognizing optimal planning frequencies ("Goldilocks zones"), such as planning every four steps rather than at every opportunity, we removed explicit reward shaping, thus allowing agents to autonomously learn optimal planning frequencies. As illustrated in Figure 9, agents trained under this strategy gradually improved their efficiency. Specifically, agents learned to execute plans less frequently (center panel) but with increased effectiveness, leading to shorter, more concise plans (right panel), and improved overall performance (left panel).
@@ D.1 Qualitative results
First, Figure 10 showcases successful human-agent collaboration, detailing how human-provided high-level plans guided the agent to a game-winning Crafter trajectory after approximately 20 attempts. Next, we demonstrate its autonomous dynamic planning. Figure 11 shows the agent interrupting an ongoing task and adaptively replanning to acquire critical food supplies when its health is low. Figure 12 further illustrates its dynamic planning capabilities, observing the agent employing multi-stage tactics: initially planning to craft a weapon, then devising a new plan to strategically position itself before engaging enemies. Finally, Figure 13 presents an execution failure, showing how an otherwise sound plan to craft an item falters because the agent fails to verify all necessary prerequisites—specifically, by not ensuring a required furnace is accessible. This is akin to the knowing-doing gap identified in BALROG [Paglieri et al., 2025a].
@@
D.2 RL Cost Penalty Ablation

### v2

@@ A.3 Dynamic Planning Prompts
Figure 4: An illustration of the agent’s input context over two timesteps, $t$ and $t+1$. At each timestep, the agent processes a chat-formatted history composed of a system prompt, user messages (green observation and gray instruction), and assistant messages (yellow plan and blue action). The agent receives the history of interactions and in this case it generates a new plan and action. In the subsequent timestep $t+1$, the input history is updated: the plan and action generated at $t$ are appended to the interaction history together with new observation, and the previous plan is removed. During experiments, to manage context length, the history provided to the agent was truncated to a maximum of 16 observations.
@@ B.1 Crafter
Crafter Hafner (2022) is an open-world survival game inspired by Minecraft. The environment is designed to benchmark reinforcement learning agents on tasks requiring generalization and long-term reasoning without extensive prior game knowledge. Crafter presents a procedurally generated 2D world where the agent must gather resources, craft tools, and defend against creatures to survive and unlock achievements. The environment provides observations as a combination of top-down pixel views of the agent’s local surroundings and non-visual data, including the agent’s health, inventory, and currently held item. The agent can perform 17 discrete actions, and its progress is measured by unlocking 22 predefined achievements, each yielding a sparse reward signal.
Our agent interacts with the Crafter environment through the BALROG Paglieri et al. (2025a) framework, which acts as a standardized wrapper. BALROG facilitates the communication between our agent, conceptualized as an underlying Large Language Model (LLM) combined with a specific prompting strategy, and the Crafter game environment. At each timestep, Crafter, via BALROG, relays the current observation to our agent. The agent, which internally maintains a history of past observations and actions, incorporates this new observation into its context. A dedicated prompt builder component within the agent updates this interaction history and formats it into a chat template. An example figure of Crafter together with current observation can be see in figure 5 on the right.
This prompt is then passed to the LLM, which processes the contextual information and generates the subsequent action as a natural language string. BALROG then translates this string into a game-compatible command for Crafter. Further details regarding the specifics of the system prompt, the wrapper’s mechanics, and the handling of action validation can be found in BALROG Paglieri et al. (2025a).
@@ C.1 Supervised Fine-Tuning
For the Supervised Fine-Tuning (SFT) phase, we generated a dataset of 1,024 trajectories using Llama 3.3 70B instruct. To ensure data diversity, we employed 16 distinct planning prompts (detailed in Appendix A), which were sampled uniformly. The planning frequency for trajectories in the dataset was also sampled uniformly from the range [2, 12]. The Llama 3.1 8B instruct model served as the baseline for all SFT experiments, with no Low-Rank Adaptation (LoRA) adapters applied during this stage (Hu et al., 2022). We used AdamW (Loshchilov & Hutter, 2019) as the optimizer throughout.
We considered two objective functions: direct action prediction and full world modeling. Our findings, presented in Figures 7 and 8, indicate that omitting world modeling yields superior results in terms of both performance and reduced catastrophic forgetting. While we initially tested world modeling as a potential regularization technique to mitigate overfitting, it ultimately proved to be a distractor for the agents. To additionally model the environment dynamics, the weights of the model had to be changed more significantly, which is reflected in the higher KL divergence, and consequently led to increased forgetting and reduced model steerability. It is worth noting that for the agent to plan dynamically during SFT, it must also model the plans themselves, exposing it to a greater number of tokens during this stage. Training was performed using the DeepSpeed ZeRO Stage 3 optimizer (Aminabadi et al., 2022) to effectively manage memory and scale operations across multiple GPUs.
@@ C.2 RLFT
During the Reinforcement Learning Fine-Tuning (RLFT) phase, we applied RSLoRA adapters (Kalajdzievski, 2023) separately to actor and critic models. Data collection involved using vLLM (Kwon et al., 2023) to host model weights and BALROG (Paglieri et al., 2025a) integrated with Ray (Moritz et al., 2018) to process outputs and handle agent-environment interactions efficiently. A custom Proximal Policy Optimization (PPO) trainer (Schulman et al., 2017) was implemented, with model weights broadcasted back to vLLM after each training epoch. The actor model was based on the 8B parameter Llama 3.1 architecture, while the critic model utilized the 1B parameter Llama 3.3 architecture (Grattafiori et al., 2024). Similar to the SFT phase, the DeepSpeed ZeRO Stage 3 optimizer was employed for memory-efficient and scalable training.
Experiments were conducted using a total of 8 GPUs, allocating 6 GPUs for training and 2 for data collection. Most experiments were executed on a node with 8xH100 GPUs and typically completed within 24–48 hours. Initially, we experimented with reward penalties targeting invalid actions, excessively long responses, and overly frequent planning. However, such explicit reward shaping often resulted in agents refraining from planning entirely. Recognizing optimal planning frequencies (”Goldilocks zones”), such as planning every four steps rather than at every opportunity, we removed explicit reward shaping, thus allowing agents to autonomously learn optimal planning frequencies. As illustrated in Figure 9, agents trained under this strategy gradually improved their efficiency. Specifically, agents learned to execute plans less frequently (center panel) but with increased effectiveness, leading to shorter, more concise plans (right panel), and improved overall performance (left panel).
@@ D.1 Qualitative results
First, Figure 11 showcases successful human-agent collaboration, detailing how human-provided high-level plans guided the agent to a game-winning Crafter trajectory after approximately 20 attempts. Next, we demonstrate its autonomous dynamic planning. Figure 12 shows the agent interrupting an ongoing task and adaptively replanning to acquire critical food supplies when its health is low. Figure 13 further illustrates its dynamic planning capabilities, observing the agent employing multi-stage tactics: initially planning to craft a weapon, then devising a new plan to strategically position itself before engaging enemies. Finally, Figure 14 presents an execution failure, showing how an otherwise sound plan to craft an item falters because the agent fails to verify all necessary prerequisites—specifically, by not ensuring a required furnace is accessible. This is akin to the knowing-doing gap identified in BALROG (Paglieri et al., 2025a).
@@
D.2 Best-of-N Upper Bound
To quantify the upper bound of performance, we conducted a Best-of-N analysis, comparing the best runs attained by each method across 100 independent runs, and 20 runs for the human collaboration. As shown in Figure 15, the SFT+RL when guided by human plans, successfully solves Crafter by mining a diamond, significantly outperforming all of the methods. This result demonstrates the impact of human steering and further validates the benefits of conditioning the models with plans.